# The extension of Schwarzschild line element to include uniformly accelerated mass

### v3

@@ 3 Metric for a uniformly accelerated Schwarzschild object
$g_{t\,t}$ $=\frac{a_{x}^{2}c^{2}t^{2}\left(2m\Sigma_{2}-\Sigma_{0}^{3/2}\right)}{\left(\Sigma_{0}^{3/2}-2\Sigma_{0}m\right)\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}}+\frac{a_{y}^{2}c^{2}t^{2}\left(2\Sigma_{1}m-\Sigma_{0}^{3/2}\right)}{\left(\Sigma_{0}^{3/2}-2\Sigma_{0}m\right)\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}}$
$-\frac{4a_{x}a_{y}mc^{2}t^{2}\Sigma_{4}}{\left(\Sigma_{0}^{3/2}-2\Sigma_{0}m\right)\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}}+\frac{c^{4}\left(\sqrt{\Sigma_{0}}-2m\right)\rho^{2}}{\sqrt{\Sigma_{0}}\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}^{2}}$ (30)
$g_{xx}$ $=\frac{1}{\Sigma_{0}^{3/2}-2\Sigma_{0}m}\Bigg{[}\frac{\sqrt{\Sigma_{0}}a_{x}^{2}c^{6}t^{2}\left(\sqrt{\Sigma_{0}}-2m\right)^{2}}{\left(-a_{xy}^{2}c^{2}t^{2}+2c^{2}(a_{x}x+a_{y}y)+(a_{x}x+a_{y}y)^{2}+c^{4}\right)^{2}}$
$+\frac{a_{x}^{2}a_{y}^{2}}{a_{xy}^{4}}\left(2\Sigma_{1}m-\Sigma_{0}^{3/2}\right)\left(\Sigma_{7}\right)^{2}+\left(2m\Sigma_{2}-\Sigma_{0}^{3/2}\right)\left\{\frac{a_{x}^{2}}{a_{xy}^{2}}\left(\Sigma_{7}\right)+1\right\}^{2}$
$-\frac{4a_{x}a_{y}m}{a_{xy}^{2}}\Sigma_{4}\left(\Sigma_{7}\right)\left\{\frac{a_{x}^{2}}{a_{xy}^{2}}\left(\Sigma_{7}\right)+1\right\}\Bigg{]}$ (31)
$g_{yy}$ $=\frac{1}{\Sigma_{0}^{3/2}-2\Sigma_{0}m}\Bigg{[}\frac{\sqrt{\Sigma_{0}}a_{y}^{2}c^{6}t^{2}\left(\sqrt{\Sigma_{0}}-2m\right)^{2}}{\left(-a_{xy}^{2}c^{2}t^{2}+2c^{2}(a_{x}x+a_{y}y)+(a_{x}x+a_{y}y)^{2}+c^{4}\right)^{2}}$
$+\frac{a_{x}^{2}a_{y}^{2}}{a_{xy}^{4}}\left(2m\Sigma_{2}-\Sigma_{0}^{3/2}\right)\left(\Sigma_{7}\right)^{2}+\left(2\Sigma_{1}m-\Sigma_{0}^{3/2}\right)\left\{\frac{a_{y}^{2}}{a_{xy}^{2}}\left(\Sigma_{7}\right)+1\right\}^{2}$
$-\frac{4a_{x}a_{y}m}{a_{xy}^{2}}\Sigma_{4}\left(\Sigma_{7}\right)\left\{\frac{a_{y}^{2}}{a_{xy}^{2}}\left(\Sigma_{7}\right)+1\right\}\Bigg{]}$ (32)
$g_{zz}$ $=\frac{1}{\Sigma_{0}^{3/2}-2\Sigma_{0}m}\Bigg{[}\frac{2m}{a_{xy}^{4}}\Bigg{[}\Big{\{}a_{xy}^{2}x-a_{x}\Big{\{}-\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}+\rho\Big{\}}\Big{\}}^{2}$
$+\Big{\{}a_{xy}^{2}y-a_{y}\Big{\{}-\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}+\rho\Big{\}}\Big{\}}^{2}\Bigg{]}-\Sigma_{0}^{3/2}\Bigg{]}$ (33)
$g_{xt}$ $=-ct\Bigg{[}\frac{a_{x}c^{4}\left(\sqrt{\Sigma_{0}}-2m\right)\rho}{\sqrt{\Sigma_{0}}\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}^{2}}+\frac{a_{x}\left(2m\Sigma_{2}-\Sigma_{0}^{3/2}\right)}{a_{xy}^{2}\left(\Sigma_{0}^{3/2}-2\Sigma_{0}m\right)\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}}$
$\times\Bigg{\{}a_{x}^{2}\left(-\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}+\rho\right)+a_{xy}^{2}\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}\Bigg{\}}$
$-\frac{2a_{x}^{2}a_{y}m\Sigma_{4}\left(-\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}+\rho\right)}{a_{xy}^{2}\left(\Sigma_{0}^{3/2}-2\Sigma_{0}m\right)\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}}$
$-\frac{2a_{y}m\Sigma_{4}\Big{\{}a_{x}^{2}\left(-\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}+\rho\right)+a_{xy}^{2}\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}\Big{\}}}{a_{xy}^{2}\left(\Sigma_{0}^{3/2}-2\Sigma_{0}m\right)\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}}$
$+\frac{a_{x}a_{y}^{2}\left(2\Sigma_{1}m-\Sigma_{0}^{3/2}\right)\left(-\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}+\rho\right)}{a_{xy}^{2}\left(\Sigma_{0}^{3/2}-2\Sigma_{0}m\right)\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}}\Bigg{]}$ (34)
$g_{yt}$ $=-ct\Bigg{[}\frac{a_{y}c^{4}\left(\sqrt{\Sigma_{0}}-2m\right)\rho}{\sqrt{\Sigma_{0}}\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}^{2}}+\frac{a_{x}^{2}a_{y}\left(2m\Sigma_{2}-\Sigma_{0}^{3/2}\right)}{a_{xy}^{2}\left(\Sigma_{0}^{3/2}-2\Sigma_{0}m\right)\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}}$
$\times\left(-\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}+\rho\right)+\frac{a_{y}\left(2\Sigma_{1}m-\Sigma_{0}^{3/2}\right)}{a_{xy}^{2}\left(\Sigma_{0}^{3/2}-2\Sigma_{0}m\right)\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}}$
$\times\left\{a_{y}^{2}\left(-\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}+\rho\right)+a_{xy}^{2}\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}\right\}$
$-\frac{2a_{x}a_{y}^{2}m\Sigma_{4}\left(-\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}+\rho\right)}{a_{xy}^{2}\left(\Sigma_{0}^{3/2}-2\Sigma_{0}m\right)\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}}-\frac{2a_{x}m\Sigma_{4}}{a_{xy}^{2}\left(\Sigma_{0}^{3/2}-2\Sigma_{0}m\right)\Big{\{}\rho^{2}-a_{xy}^{2}c^{2}t^{2}\Big{\}}}$
$\times\Bigg{\{}a_{y}^{2}\left(-\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}+\rho\right)+a_{xy}^{2}\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}\Bigg{\}}\Bigg{]}$ (35)
$g_{zt}$ $=\frac{2mct\Sigma_{3}z}{\Sigma_{0}\left(\sqrt{\Sigma_{0}}-2m\right)\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}}$ (36)
$g_{xy}$ $=\frac{1}{\Sigma_{0}^{3/2}-2\Sigma_{0}m}\Bigg{[}\frac{\sqrt{\Sigma_{0}}a_{x}a_{y}c^{6}t^{2}\left(\sqrt{\Sigma_{0}}-2m\right)^{2}}{\left(-a_{xy}^{2}c^{2}t^{2}+2c^{2}(a_{x}x+a_{y}y)+(a_{x}x+a_{y}y)^{2}+c^{4}\right)^{2}}$
$+\frac{a_{x}a_{y}}{a_{xy}^{2}}\left(2m\Sigma_{2}-\Sigma_{0}^{3/2}\right)\left\{\frac{a_{x}^{2}}{a_{xy}^{2}}\left(\Sigma_{7}\right)+1\right\}\left(\Sigma_{7}\right)+\frac{a_{x}a_{y}}{a_{xy}^{2}}\left(2\Sigma_{1}m-\Sigma_{0}^{3/2}\right)$
$\times\left\{\frac{a_{y}^{2}}{a_{xy}^{2}}\left(\Sigma_{7}\right)+1\right\}\left(\Sigma_{7}\right)-2m\Sigma_{4}\left\{\frac{a_{x}^{2}}{a_{xy}^{2}}\left(\Sigma_{7}\right)+1\right\}\left\{\frac{a_{y}^{2}}{a_{xy}^{2}}\left(\Sigma_{7}\right)+1\right\}$
$-\frac{2a_{x}^{2}a_{y}^{2}m\Sigma_{4}\left(-\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}+\rho\right)^{2}}{a_{xy}^{4}\left(-a_{xy}^{2}c^{2}t^{2}+2c^{2}(a_{x}x+a_{y}y)+(a_{x}x+a_{y}y)^{2}+c^{4}\right)}\Bigg{]}$ (37)
$g_{yz}$ $=-\frac{2m\Sigma_{6}z}{a_{xy}^{2}\Sigma_{0}\left(\sqrt{\Sigma_{0}}-2m\right)\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}}$ (38)
$g_{xz}$ $=-\frac{2m\Sigma_{5}z}{a_{xy}^{2}\Sigma_{0}\left(\sqrt{\Sigma_{0}}-2m\right)\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}}$ (39)
in which the notations $\Sigma_{0}$, $\Sigma_{1}$, $\Sigma_{2}$, $\Sigma_{3}$, $\Sigma_{4}$, $\Sigma_{5}$, $\Sigma_{6}$, $\Sigma_{7}$, $\rho_{0}$ and $\rho$ are used to represent as follows:
$\Sigma_{0}$ $=\left\{a_{x}\left(\rho_{0}\right)+x\right\}^{2}+\left\{a_{y}\left(\rho_{0}\right)+y\right\}^{2}+z^{2}$ (40)
$\Sigma_{1}$ $=\left\{a_{x}\left(\rho_{0}\right)+x\right\}^{2}+z^{2}$ (41)
$\Sigma_{2}$ $=\left\{a_{y}\left(\rho_{0}\right)+y\right\}^{2}+z^{2}$ (42)
$\Sigma_{3}$ $=a_{x}\left\{a_{x}\left(\rho_{0}\right)+x\right\}+a_{y}\left\{a_{y}\left(\rho_{0}\right)+y\right\}$ (43)
$\Sigma_{4}$ $=\left\{a_{x}\left(\rho_{0}\right)+x\right\}\left\{a_{y}\left(\rho_{0}\right)+y\right\}$ (44)
$\Sigma_{5}$ $=a_{x}^{3}x\left\{a_{x}\left(\rho_{0}\right)+x\right\}+a_{x}^{2}c^{2}\left\{a_{x}\left(\rho_{0}\right)+x\right\}+a_{x}^{2}a_{y}y\left\{a_{x}\left(\rho_{0}\right)+x\right\}$
$+a_{x}^{2}a_{y}x\left\{a_{y}\left(\rho_{0}\right)+y\right\}+a_{x}a_{y}^{2}y\left\{a_{y}\left(\rho_{0}\right)+y\right\}+a_{x}a_{y}c^{2}\left\{a_{y}\left(\rho_{0}\right)+y\right\}$
$-a_{x}^{2}\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}\left\{a_{x}\left(\rho_{0}\right)+x\right\}+a_{xy}^{2}\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}$
$\times\left\{a_{x}\left(\rho_{0}\right)+x\right\}-a_{x}a_{y}\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}\left\{a_{y}\left(\rho_{0}\right)+y\right\}$ (45)
$\Sigma_{6}$ $=a_{x}^{2}a_{y}x\left\{a_{x}\left(\rho_{0}\right)+x\right\}+a_{y}^{3}y\left\{a_{y}\left(\rho_{0}\right)+y\right\}+a_{x}a_{y}^{2}y\left\{a_{x}\left(\rho_{0}\right)+x\right\}$
$+a_{y}^{2}c^{2}\left\{a_{y}\left(\rho_{0}\right)+y\right)\}+a_{x}a_{y}^{2}x\left\{a_{y}\left(\rho_{0}\right)+y\right\}+a_{x}a_{y}c^{2}\left\{a_{x}\left(\rho_{0}\right)+x\right\}$
$-a_{y}^{2}\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}\left\{a_{y}\left(\rho_{0}\right)+y\right\}+a_{xy}^{2}\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}$
$\times\left\{a_{y}\left(\rho_{0}\right)+y\right\}-a_{x}a_{y}\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}$
$\times\left\{a_{x}\left(\rho_{0}\right)+x\right\}$ (46)
$\Sigma_{7}$ $=\frac{\rho}{\sqrt{\rho^{2}-a_{xy}^{2}c^{2}t^{2}}}-1$ (47)
$\rho_{0}$ $=\sqrt{\left(\frac{\rho}{a_{xy}^{2}}\right)^{2}-\frac{c^{2}t^{2}}{a_{xy}^{2}}}-\frac{\rho}{a_{xy}^{2}}$ (48)
$\text{and\,\,\,\,}\rho$ $=a_{x}x+a_{y}y+c^{2}$ (49)
@@ 4 Discussion of results and conclusions
Most importantly, our tools will be useful in the studies of gravitational waves. The gravitational-wave signal observed by a detector in the vicinity of Earth is actually approximated by the waves along some geodesic related to distance from the source. Thus, if we estimate a quantity as detected by the Gravitational Wave (GW) detector, we approximate the signal an inertial observer measures as a function of proper time [28]. Furthermore, the Rindler framework can be used to study the behavior of detectors in curved spacetime, like those around black holes, where the spacetime curvature is significant. In the context of GW detection, the Rindler transformation can be used to analyze the interaction between GWs and detectors, particularly when considering the quantum nature of both. The Rindler transformation provides a valuable tool for understanding the perspectives of accelerated observers and the thermal effects associated with horizons in both flat and curved spacetime. This, in turn, helps in exploring complex phenomena like the behavior of GW detectors in the presence of strong gravitational fields and the interplay between gravitational waves, quantum fields, and the Unruh effect. The Rindler transformation helps us understand how GWs can transfer energy to the internal states of a detector, potentially affecting its coherence and response [29]. By studying the response of detectors in the Rindler frame, researchers can gain insights into the fundamental nature of quantum fields in curved spacetime and the properties of GWs themselves [30]. The Rindler transformation in the background of curved spacetime, provides a valuable theoretical tool for bridging the gap between general relativity, quantum field theory, and the experimental realm of GW detection [31]. Thus our present study on how Lorentz frame / Rindler frame behaves in the background of curved spacetime (in our case Schwarzschild field) will have relevance in the research on gravitational wave and its detection .
@@
Appendix A Metrics for particles having velocity close to the speed of light ($c$)
The Aichelburg and Sexl [5] derived a metric for massless particles from the isotropic form of the Schwarzschild metric using a singular Lorentz transformation, which can be reproduced as follows  [^3] In original derivation Aichelburg and Sexl [5] considered speed of light $c=1$.:
$ds^{2}$ $=c^{2}dt^{\prime 2}-d\chi^{\prime 2}-d\lambda^{\prime 2}-d\sigma^{\prime 2}-\frac{4GE\Big{(}cdt^{\prime}-d\chi^{\prime}\Big{)}^{2}}{c^{4}\left(\chi^{\prime}-ct^{\prime}\right)}$ (A.1)
Now in order to examine the metric (10) in the limit of $v_{x}\longrightarrow c$ we proceed as follows: Putting $W=\sqrt{\mathcal{R}}$ in the metric (10) and expanding the metric components in the power of $2m/W$ we can obtain as follows:
$g_{tt}$ $\approx 1+\gamma^{2}\Bigg{[}-1-\frac{v_{x}^{2}}{c^{2}}+\frac{v_{x}^{2}}{c^{2}}\cdot\frac{\left(y^{2}+z^{2}\right)}{W}\Bigg{]}\left(\frac{2m}{W}\right)+$... (A.2)
$g_{xx}$ $\approx-1+\gamma^{2}\Bigg{[}-1-\frac{v_{x}^{2}}{c^{2}}+\frac{\left(y^{2}+z^{2}\right)}{W}\Bigg{]}\left(\frac{2m}{W}\right)+$... (A.3)
$g_{yy}$ $\approx-1+\Bigg{[}-1+\frac{\gamma^{2}(x-v_{x}t)^{2}+z^{2}}{W}\Bigg{]}\left(\frac{2m}{W}\right)+$... (A.4)
$g_{zz}$ $\approx-1+\Bigg{[}-1+\frac{\gamma^{2}(x-v_{x}t)^{2}+y^{2}}{W}\Bigg{]}\left(\frac{2m}{W}\right)+$... (A.5)
$g_{tx}$ $\approx\frac{\gamma^{2}v_{x}^{2}}{c^{2}W^{2}}\Big{\{}2\gamma^{2}(x-v_{x}t)^{2}+y^{2}+z^{2}\Big{\}}\left(\frac{2m}{W}\right)+$... (A.6)
$g_{ty}$ $\approx\gamma^{2}\frac{v_{x}^{2}}{c^{2}W^{2}}y(x-v_{x}t)\left(\frac{2m}{W}\right)+$... (A.7)
$g_{tz}$ $\approx\gamma^{2}\frac{v_{x}^{2}}{c^{2}W^{2}}z(x-v_{x}t)\left(\frac{2m}{W}\right)+$... (A.8)
$g_{xy}$ $\approx-\frac{\gamma^{2}}{W^{2}}y(x-v_{x}t)\left(\frac{2m}{W}\right)+$... (A.9)
$g_{xz}$ $\approx-\frac{\gamma^{2}}{W^{2}}z(x-v_{x}t)\left(\frac{2m}{W}\right)+$... (A.10)
$g_{yz}$ $\approx-\frac{yz}{W^{2}}\left(\frac{2m}{W}\right)+$... (A.11)
Now using the relativistic mass $M_{r}$ and energy $E$ relation: $E=M_{r}c^{2}$, we write $E=\frac{Mc^{2}}{\sqrt{1-\frac{v_{x}^{2}}{c^{2}}}}$ and hence $m=\frac{GM}{c^{2}}=\frac{GE}{c^{4}}\sqrt{1-\frac{v_{x}^{2}}{c^{2}}}$. Then, substituting $m=\frac{GE}{c^{4}}\sqrt{1-\frac{v_{x}^{2}}{c^{2}}}$ in the metric components (A.2)–(A.11) and applying the limit $v_{x}\rightarrow c$ we obtain as:
$\lim_{v_{x}\to c}g_{tt}$ $=1-\left\{\frac{4GE}{c^{4}(x-ct)}\right\},\hskip 28.45274pt\lim_{v_{x}\to c}g_{xx}=-1-\left\{\frac{4GE}{c^{4}(x-ct)}\right\}$ (A.12)
$\lim_{v_{x}\to c}g_{yy}$ $=-1,\hskip 28.45274pt\lim_{v_{x}\to c}g_{zz}=-1,\hskip 28.45274pt\lim_{v_{x}\to c}g_{tx}=\left\{\frac{4GE}{c^{4}(x-ct)}\right\}$
$\lim_{v_{x}\to c}g_{ty}$ $=0,\hskip 17.07182pt\lim_{v_{x}\to c}g_{tz}=0,\hskip 17.07182pt\lim_{v_{x}\to c}g_{xy}=0,\hskip 17.07182pt\lim_{v_{x}\to c}g_{xz}=0,\hskip 17.07182pt\lim_{v_{x}\to c}g_{yz}=0$
Therefore the spacetime metric (10) at $v_{x}\rightarrow c$ can be written from (A.12) as follows:
$ds^{2}$ $=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}-\frac{4GE}{c^{4}(x-ct)}\Big{\{}cdt-x\Big{\}}^{2}$ (A.13)
Appendix B Transformation relations for uniformly accelerated frames
B.1 Inverse Rindler transformation

### v4

@@ 4 Metric for a uniformly accelerated Schwarzschild object
$g_{t\,t}$ $=\frac{c^{4}}{\left(\xi_{2}\right)^{2}}\left(1-\frac{2m}{\Delta_{3}-m}\right)\left(a_{x}x+a_{y}y+c^{2}\right)^{2}$
$\hskip 56.9055pt-\frac{c^{2}t^{2}}{\xi_{2}}\left(1+\frac{2m}{\Delta_{3}-m}\right)\Big\{a_{x}^{2}+a_{y}^{2}\Big\}$ (36)
$g_{xx}$ $=-\left(1+\frac{2m}{\Delta_{3}-m}\right)\Bigg[\left(\frac{a_{x}^{2}\xi_{1}}{a^{2}}+1\right)^{2}+\frac{a_{x}^{2}a_{y}^{2}\xi_{1}^{2}}{a^{4}}\Bigg]$
$\hskip 113.81102pt+\frac{a_{x}^{2}c^{6}t^{2}}{\left(\xi_{2}\right)^{2}}\left(1-\frac{2m}{\Delta_{3}-m}\right)$ (37)
$g_{yy}$ $=-\left(1+\frac{2m}{\Delta_{3}-m}\right)\Bigg[\left(\frac{a_{y}^{2}\xi_{1}}{a^{2}}+1\right)^{2}+\frac{a_{x}^{2}a_{y}^{2}\xi_{1}^{2}}{a^{4}}\Bigg]$
$\hskip 56.9055pt+\frac{a_{y}^{2}c^{6}t^{2}}{\left(\xi_{2}\right)^{2}}\left(1-\frac{2m}{\Delta_{3}-m}\right)$ (38)
$g_{zz}$ $=-\Bigg\{1+\frac{2m}{\Delta_{3}-m}\Bigg\}$ (39)
$g_{xt}$ $=a_{x}ct\Bigg[\frac{1}{\sqrt{\xi_{2}}}\Bigg\{\frac{\xi_{1}}{a^{2}}\Big(a_{x}^{2}+a_{y}^{2}\Big)+1\Bigg\}\left(1+\frac{2m}{\Delta_{3}-m}\right)$
$-\frac{c^{5}}{\left(\xi_{2}\right)^{2}}\left(1-\frac{2m}{\Delta_{3}-m}\right)\left(a_{x}x+a_{y}y+c^{2}\right)\Bigg]$ (40)
$g_{yt}$ $=a_{y}ct\Bigg[\frac{1}{\sqrt{\xi_{2}}}\Bigg\{\frac{\xi_{1}}{a^{2}}\Big(a_{x}^{2}+a_{y}^{2}\Big)+1\Bigg\}\left(1+\frac{2m}{\Delta_{3}-m}\right)$
$-\frac{c^{5}}{\left(\xi_{2}\right)^{2}}\left(1-\frac{2m}{\Delta_{3}-m}\right)\left(a_{x}x+a_{y}y+c^{2}\right)\Bigg]$ (41)
$g_{xy}$ $=a_{x}a_{y}\Bigg[\frac{c^{6}t^{2}}{\left(\xi_{2}\right)^{2}}\left(1-\frac{2m}{\Delta_{3}-m}\right)-\left(1+\frac{2m}{\Delta_{3}-m}\right)$
$\hskip 56.9055pt\times\frac{\xi_{1}}{a^{2}}\Bigg\{\frac{\xi_{1}}{a^{2}}\left(a_{x}^{2}+a_{y}^{2}\right)+2\Bigg\}\Bigg]$ (42)
in which $\Delta_{3}$, $\xi_{1}$ and $\xi_{2}$ are used to represent as follows:
$\Delta_{3}$ $=\Bigg\{\Big\{x+\frac{a_{x}}{a^{2}}\Big(\sqrt{\xi_{2}}-a_{x}x-a_{y}y-c^{2}\Big)\Big\}^{2}$
$+\Big\{y+\frac{a_{y}}{a^{2}}\Big(\sqrt{\xi_{2}}-a_{x}x-a_{y}y-c^{2}\Big)\Big\}^{2}+z^{2}\Bigg\}^{1/2}$ (43)
$\xi_{1}$ $=\left\{\frac{1}{\sqrt{\xi_{2}}}\Big(a_{x}x+a_{y}y+c^{2}\Big)-1\right\}$ (44)
$\text{and\,\,\,\,}\xi_{2}$ $=\left(a_{x}x+a_{y}y+c^{2}\right)^{2}-a^{2}c^{2}t^{2}$ (45)
@@ 5 Discussion of results and conclusions
Most importantly, our tools will be useful in the studies of gravitational waves. The gravitational-wave signal observed by a detector in the vicinity of Earth is actually approximated by the waves along some geodesic related to distance from the source. Thus, if we estimate a quantity as detected by the Gravitational Wave (GW) detector, we approximate the signal an inertial observer measures as a function of proper time [32]. Furthermore, the Rindler framework can be used to study the behavior of detectors in curved spacetimes, like those around black holes, where the spacetime curvature is significant. In the context of GW detection, the Rindler transformation can be used to analyze the interaction between GWs and detectors, particularly when considering the quantum nature of both. The Rindler transformation provides a valuable tool for understanding the perspectives of accelerated observers and the thermal effects associated with horizons in both flat and curved spacetime. This, in turn, helps in exploring complex phenomena like the behavior of GW detectors in the presence of strong gravitational fields and the interplay between gravitational waves, quantum fields, and the Unruh effect. The Rindler transformation helps us understand how GWs can transfer energy to the internal states of a detector, potentially affecting its coherence and response [33]. By studying the response of detectors in the Rindler frame, researchers can gain insights into the fundamental nature of quantum fields in curved spacetime and the properties of GWs themselves [34]. The Rindler transformation in the background of curved spacetime, provides a valuable theoretical tool for bridging the gap between general relativity, quantum field theory, and the experimental realm of GW detection [35]. Thus our present study on how Lorentz frame / Rindler frame behaves in the background of curved spacetime (in our case Schwarzschild field) will have relevance in the research on gravitational wave and its detection .
@@
Appendix A Transformation relations for uniformly accelerated frames
A.1 Inverse Rindler transformation